AGw Commissioning with IRTC on 18-19 October 2008

Observers: J. Hill on mountain and A. Rakich in Tucson plus J. Schmidt working on DIMM
Software support: C. Biddick in Tucson (first two hours)
Telescope Operator: A. Ceranski

Summary

Seeing variable between 0.6 and 1.8 arcsec. Up to 2 mags of clouds, but clear most of night. Winds decreasing down to almost nothing.

Took many cubes of IRTC data for vibration/jitter analysis.

Confirmed that collimating off-axis, with AGw X = 0 and X non-zero, every image we took through the night had good correction on the IRTC (apart from some suffering from a known issue with GCS). Curvature sensing data taken to prove it.

Got good data on wavefront sensor scale factors towards end of night, problems with SF for Z4 and especially Z11.

Details

Times reported in UT below

IRTC data in /Repository/20081019_IRTC/

Initial setup

01:10 Open telescope with partly cloudy skies

01:48 pointto M5_0707 for initial pointing

IRTC is near saturation in 0.5 sec exposures. Something is wrong with TEC as temperature of camera is reporting 283K.

02:10 spiral search looking for star on IRTC. Why are 30 arcsec offsets taking 25 sec?

Issue 1722 again

02:21 pointto M5_0707 ACQUIRE gs=0

Oct 18 19:22:06 lbtmu107 LBT_left_GCS: AGWUnit initialized, calling init() ...
Oct 18 19:22:08 lbtmu107 LBT_left_GCS: AGWunit init(): _filterPosition or _focus read back from AGW server is negative indicating an error.
Oct 18 19:22:08 lbtmu107 LBT_left_GCS: init() returned an error, indicating the AGW unit is not accessible.

02:24 restarted check4nport and oacserverlog - didn't help

02:26 restarted GCSL - didn't help

02:28 stopAGW / startAGW

02:32 The problem was that the probe was parked at -0.001 mm.

02:38 readGuideCam and we see two faint stars that are approximately collimated.

Spiraling in DETXY REL rather than RADEC ABS

02:40 cl< spiral.cl Now offsets are taking 4 sec. ?? But then increase to 25 sec. ??

Oops we were offsetting in DETXY and not RADEC. Does REL work in DETXY mode?

Actually acquiring the star

03:00 cl < spiral.cl Found star at -60 50. IE=-27 CA-27.

03:13 offset 0 10 moves star down on Guider image.

03:14 pointto M5_0707 TRACK and now we happily see it on IRTC - collimation model is quite good now since last night's updates based on temperature equilibirum.

03:19 pointto M5_0738 TRACK

03:21 pointto M5_0738 gs=4 ACTIVE, but couldn't find star

03:24 pointto M5_0738 gs=0 ACTIVE, found star but too bright of course

03:29 pointto M5_0738 gs=3 ACTIVE

Oct 18 20:29:11 lbtmu107 LBT_left_GCS: full CCD readout requested, now waiting for image to arrive.
Oct 18 20:29:23 lbtmu107 LBT_left_GCS: 1. reception of image failed to arrive within timeout period of 7659 ms, trying again in 8 seconds...
Oct 18 20:29:23 lbtmu107 LBT_left_GCS: full CCD readout requested, now waiting for image to arrive.
Oct 18 20:29:35 lbtmu107 LBT_left_GCS: 2. reception of image failed to arrive within timeout period of 7659 ms, trying again in 8 seconds...
Oct 18 20:29:35 lbtmu107 LBT_left_GCS: full CCD readout requested, now waiting for image to arrive.
Oct 18 20:29:47 lbtmu107 LBT_left_GCS: 3. reception of image failed to arrive within timeout period of 7659 ms, trying again in 8 seconds...
Oct 18 20:29:47 lbtmu107 LBT_left_GCS: full CCD readout requested, now waiting for image to arrive.
Oct 18 20:29:58 lbtmu107 LBT_left_GCS: 4. reception of image failed to arrive within timeout period of 7659 ms, trying again in 8 seconds...
Oct 18 20:29:58 lbtmu107 LBT_left_GCS: full CCD readout requested, now waiting for image to arrive.
Oct 18 20:30:10 lbtmu107 LBT_left_GCS: fatal error: 5. reception of image failed to arrive within timeout period of 7659 ms, giving up.
Oct 18 20:30:21 lbtmu107 LBT_Networkserver: GCSL died unexpectedly
Oct 18 20:30:21 lbtmu107 LBT_Networkserver: GCSL is not active
Oct 18 20:30:21 lbtmu107 LBT_Networkserver: dropping GCSL

03:31 Restart GCSL

03:31 pointto M5_0738 gs=3 ACTIVE , several faint stars

03:33 pointto M5_0738 gs=2 ACTIVE

03:37 Refocus and pointto again, but we still don't find an off-axis star of sensible brightness. Is something amiss with the probe coordinate system for AGw while authorized as IRTC? pointto and PCS seem to agree on the XY coordinates. It must be that these guidestars are too faint.

Johannes reports that DIMM is measuring 0.8 arcsec seeing. Polaris monitor reports 2.5 arcsec - apparently contaminated by windshake.

03:40 pointto BS9188 gs=0 ACTIVE, we are collimating without problems on-axis.

irs problem

03:52 pointto BS9188 gs=29 ACTIVE doesn't do it, but reports success without an error message?

IRTC> !irc SetHotspot 159.22 128.58
irc: Network.cpp:843: Ice::DNSException:
DNS error: Name or service not known
host: irs
Exit 1

03:57 restart irs and that cures the problem.

Looking for residual astigmatism after we are collimated

04:04 pointto M5_0855 gs=1 ACTIVE

Star is at 156 154 on IRTC (measured by region of interest) but depends on the rotator angle

04:12 stopWFS from GUI, send 1000 nm Z4 - shows now obvious elongation on IRTC, but we see the elongation when we send 1000 nm of Z6 to M1.

04:04 pointto M5_0855 gs=6 ACTIVE - no aberration obvious above the seeing limit.

04:23 pointto M5_0855 gs=5 ACTIVE at PA=20.

04:27 Bad astigmatism applied by GCS WFS, got up to 3 arcsec long

Johannes reports 1.6 - 1.7 arcsec seeing - confirming what we see on IRTC

04:23 pointto M5_0855 gs=1 ACTIVE at PA=20. (89 mm off-axis) WFS seems to have dialed out the astigmatism.

04:35 pointto M5_0855 gs=2 ACTIVE at PA=201.69 missed star

04:36 pointto M5_0855 TRACK PAM- IE=-40 CA=-29

04:39 pointto M5_0855 gs=2 ACTIVE at PA=201.69 below threshold

04:41 put GCS countThreshold back to 100 (we had it up at 300 last night)

Testing Guiding and WFSing in partly cloudy weather. (plus practice collecting pupils)

04:48 We've now entered the phase of testing guiding in 1-2 mags of clouds.

04:54 pointto M5_0855 gs=2 ACQUIRE adjust IE to center pupil on IRTC (not real pointing)

DIMM says 1.2

Andrew moving +0.8 mm then -0.8 mm on M2 to make intra/extrafocal pupils.

cube 0001 +0.8 sequence of 30
cube 0002 -0.8 sequence of 30

cube 0003 -0.8 sequence of 80
cube 0004 +0.8 sequence of 80 -damaged by clouds, so abandon EF pupils by now.

05:15 problem with takepic 0.1

IRTC> !irc GetFilter
0 H 1.6

05:15 pointto M5_0855 gs=2 Tel="ACTIVE"

05:19 takepic 0.1 -> irtc.20081019.051923 but image is only readable by root.

05:22 takepic 0.02

05:44 100x10msec cube 0005

05:46 1000x10 msec cube 0006 while guiding and WFSing.

05:53 Aberrations on the guidestar explode to 700 nm rms just as Andrew returns with his coffee.

05:58 Interesting 5-pointed stars. I've got to show these to George Rieke!

06:02 Aberrations have residual astigmatism.

Johannes reports seeing is 1.2 arsec.

06:06 Now the star is looking round again.

cube 0009 and 0010

06:20pointto M5_0855 gs=2 Tel="ACQUIRE"

06:21 +0.8 mm cube 0011, -0.8 cube 0012

takepic 0.4 irtc.20081019.062333.fit

06:25 +0.8 mm cube 0013, -0.8 mm cube 0014

irtc.20081019.062727.fits

06:29 pointto M5_0855 gs=6 Tel="ACTIVE" bright, so WFS using 15.4 sec.

06:37 notice image bouncing, pauseGuiding, 1000x20ms cube 0015, 0016

06:40 resumeGuiding

06:49 Brief experiment with starting and stopping M1 servo loops. No striking effect, but seeing not very good.

06:50 resumeGuiding, startWFS

06:53 pauseGuiding

06:55 +0.8 80x400 ms cube 0017

06:56 -0.8 80x400ms cube 0018

irtc.20081019.065638.fits

06:58 resumeGuiding

07:00 pauseGuiding Sent to PSF 23:59:42

07:01 +0.8 mm 80x400msec cube 0019

07:03 -0.8 mm 80x400msec cube 0020

Oct 18 23:59:41 lbtmu107 LBT_left_GCS: WFS: z(22) = 0.0000
Oct 18 23:59:42 lbtmu107 LBT_left_GCS: WFS:writing description for WFS image to: /tmp/wfscimage000643.info
Oct 18 23:59:42 lbtmu107 LBT_left_GCS: PROFILING: WFS: sending Zernikes to PSF...
Oct 18 23:59:42 lbtmu107 LBT_left_GCS: interrupted while waiting for the PSF to apply Zernikes (stopWFSing or pauseWFSing)
Oct 18 23:59:43 lbtmu107 LBT_left_GCS: WFS Cycle Time: 44.984085   WF Cam Exposure Time: 15000
Oct 18 23:59:43 lbtmu107 LBT_left_GCS: WFS Cycle Time: 44.984085   WF Cam Exposure Time: 15000
Oct 18 23:59:52 lbtmu107 LBT_left_GCS: WFS thread stopped.

Oct 18 23:59:41 lbtmu101 LBT_PSFL: PSFL:PSFSecondaryMirror:setGlobalOffsets executed
Oct 18 23:59:41 lbtmu101 LBT_PSFL: ACTOPT: Secondary collimated
Oct 18 23:59:42 lbtmu101 LBT_PSFL: PSFL:setZernikes executed
Oct 18 23:59:42 lbtmu101 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 -79.1562 102.457 7.77238 6.05995 -149.692 6.39079 -11.2959 22.5685 0 0 0 0 0 0 0 0 0 0 0
Oct 18 23:59:43 lbtmu101 LBT_PSFL: ACTOPT: Primary forces changed
Oct 18 23:59:45 lbtmu101 LBT_PSFL: ACTOPT: Primary collimated

07:04 resumeGuiding

Jitter Test for M2 Hexapod Servos

07:12 'K'illed leg #6 of M2 hexapod servo from OSS M2 GUI.

07:19 all legs of M2 hexapod now open loop with #1K, #2K, etc.

07:21 Block guiding corrections on PCS, stopWFS

07:22 80x40msec@10Hz cube 0021

07:23 1000x40msec@10Hz cube 0022, 0023 also with guider image data but not guiding.

07:28 reenable guiding in PCS.

07:30 home, init and recollimate M2

07:33 pointto M5_0855 gs=6 Tel="ACTIVE"

Total +1.055 -1.146 1.397 -18.25 -13.35 0 EL=37.4 T=6.86
Table +2.756 +2.174 +0.126 +39.42 -14.92
Temp -0.548 -0.959 +0.685 0 0
Active -0.366 -0.370 -0.074 -3.98 +3.92
Global -0.800 -2.000 +0.670 -54 -2

Oct 19 00:42:50 lbtmu107 LBT_left_GCS: PROFILING: requesting image...
Oct 19 00:42:51 lbtmu107 LBT_left_GCS: error requesting an image in GuidingThread. Lost connection?

07:45 pointto M5_0855 gs=6 Tel="ACTIVE"

Oct 19 00:46:15 lbtmu107 LBT_left_GCS: GCSL:startGuiding executed
Oct 19 00:46:15 lbtmu107 LBT_left_GCS: PROFILING: autoAcquireGuidestar called
Oct 19 00:46:15 lbtmu107 LBT_left_GCS: PROFILING: checking telescope and probe on target ...
Oct 19 00:46:16 lbtmu107 LBT_left_GCS: PROFILING: telescope and probe on target.
Oct 19 00:46:16 lbtmu107 LBT_left_GCS: telescope and AGw probe reached target positions. SFP x,y: -0.096377, 187.334564
Oct 19 00:46:16 lbtmu107 LBT_left_GCS: PROFILING: requesting full readout ...
Oct 19 00:46:16 lbtmu107 LBT_left_GCS: exposure time for full readout is 500 ms
Oct 19 00:46:16 lbtmu107 LBT_left_GCS: full CCD readout requested, now waiting for image to arrive.
Oct 19 00:46:24 lbtmu107 LBT_left_GCS: 1. reception of image failed to arrive within timeout period of 6159 ms, trying again in 7 seconds...
Oct 19 00:46:24 lbtmu107 LBT_left_GCS: full CCD readout requested, now waiting for image to arrive.
Oct 19 00:46:32 lbtmu107 LBT_left_GCS: 2. reception of image failed to arrive within timeout period of 6159 ms, trying again in 7 seconds...
Oct 19 00:46:32 lbtmu107 LBT_left_GCS: full CCD readout requested, now waiting for image to arrive.
Oct 19 00:46:40 lbtmu107 LBT_left_GCS: 3. reception of image failed to arrive within timeout period of 6159 ms, trying again in 7 seconds...
Oct 19 00:46:40 lbtmu107 LBT_left_GCS: full CCD readout requested, now waiting for image to arrive.
Oct 19 00:46:48 lbtmu107 LBT_left_GCS: 4. reception of image failed to arrive within timeout period of 6159 ms, trying again in 7 seconds...
Oct 19 00:46:48 lbtmu107 LBT_left_GCS: full CCD readout requested, now waiting for image to arrive.
Oct 19 00:46:56 lbtmu107 LBT_left_GCS: fatal error: 5. reception of image failed to arrive within timeout period of 6159 ms, giving up.
Oct 19 00:47:06 lbtmu107 LBT_Networkserver: GCSL died unexpectedly
Oct 19 00:47:06 lbtmu107 LBT_Networkserver: GCSL is not active
Oct 19 00:47:06 lbtmu107 LBT_Networkserver: dropping GCSL

07:47 Restart GCSL

07:48 pointto M5_0855 gs=6 Tel="ACTIVE", but this star is getting fairly low.

Data to evaluate the collimation with probe at non-zero X

07:50 pointto M5_0094 gs=2 Tel="ACTIVE" at EL=83.6 (too high?) GCS grabbed onto an impressively defocussed star. But the acquisition failed. Star moved ~ 1 arcminute out of field before drifting back during the acquisition offset.

Oct 19 00:53:34 lbtmu105 irs: [I] adjusted exposure time to 4600 ms
Oct 19 00:53:34 lbtmu105 irs: [I] timed out after 10 seconds waiting for the PCS to accept initial guide correction

07:56 Zglobal =+0.47 to tweak focus

07:58 pointto M5_0094 gs=2 Tel="TRACK" PAM- IE=-43 CA=-32

08:02 pointto M5_0094 gs=2 Tel="ACTIVE" @PA=244.68 X=0 Y=-74.38 mm

08:10 pauseGuiding

08:13 -0.8 mm M1 80x500msec@10Hz cube 0024 but bkgd is high compared to star

08:16 +0.8 mm M1 cube 0025

./irtc.20081019.081910.fits

08:24 pointto M5_0094 gs=11 Tel="ACTIVE" PA=224.68 X=122.472 Y= -135.198

08:33 Aaron blocks guiding in PCS, but this doesn't stop WFSing except when we defocus the star.

08:34 -0.8 mm 80x500msec@10Hz cube 0026

08:xx +0.8 mm 80x500msec@10Hz cube 0027

irtc.20081019.084456.fits

08:48 pointto M5_0182 gs=2 Tel="ACTIVE"

08:51 pointto M5_0182 gs=2 Tel="TRACK" PAM- -> IE=-35 CA=-33

08:54 pointto M5_0182 gs=2 Tel="ACTIVE" - fails - couldn't identify star in acquire000348.

Oct 19 01:54:20 lbtmu107 LBT_left_GCS: PROFILING: searching for star in full CCD readout...
Oct 19 01:54:20 lbtmu107 LBT_left_GCS: windowsize around hotspot is 80.
Oct 19 01:54:20 lbtmu107 LBT_left_GCS: subtracting background noise for WWPA guide image: 1327
Oct 19 01:54:20 lbtmu107 LBT_left_GCS: WWPA centroid close to image center, verifying star does exist...starpix = 0.00
Oct 19 01:54:20 lbtmu107 LBT_left_GCS: star was not bright enough in center(287,259) to be accepted: count = 0, threshold = 0
Oct 19 01:54:20 lbtmu107 LBT_left_GCS: windowsize around hotspot is 120.
Oct 19 01:54:20 lbtmu107 LBT_left_GCS: subtracting background noise for WWPA guide image: 1327
Oct 19 01:54:20 lbtmu107 LBT_left_GCS: star found at 279.762909, 274.703033. fine tuning with windowsize = 120
Oct 19 01:54:20 lbtmu107 LBT_left_GCS: subtracting background noise for WWPA guide image: 1327
Oct 19 01:54:20 lbtmu107 LBT_left_GCS: star fine tuned at 267.182953, 295.420471. tuning step now with windowsize = 100
Oct 19 01:54:20 lbtmu107 LBT_left_GCS: subtracting background noise for WWPA guide image: 1327
Oct 19 01:54:21 lbtmu107 LBT_left_GCS: star fine tuned at 255.024078, 310.151703. last tuning step now with windowsize = 80
Oct 19 01:54:21 lbtmu107 LBT_left_GCS: subtracting background noise for WWPA guide image: 1327
Oct 19 01:54:21 lbtmu107 LBT_left_GCS: PROFILING: end of search.
Oct 19 01:54:21 lbtmu107 LBT_left_GCS: found guide star in image at 246.721985,318.783264.
Oct 19 01:54:21 lbtmu107 LBT_left_GCS: peak value is less or equal zero!

08:59 IE=-29 CA=-37

09:00 pointto M5_0182 gs=2 Tel="ACTIVE" succeeds @PA=119.48

09:02 pointto M5_0182 gs=9 Tel="ACTIVE" PA=119.48

09:05 pointto M5_0182 gs=7 Tel="ACTIVE" PA=119.48 X=112.458 Y=94.296

09:10 pauseGuiding

09:11 +0.8 mm M2 80x100msec@10Hz cube 0028

09:12 -0.8 mm M2 80x100msec@10Hz cube 0029

09:14 0.0 1000x5msec@50Hz cube 0030

09:16 0.0 3000x5msec@20Hz cube 0031

irtc.20081019.092140.fits

09:28 star starts a strong vibration, so start new cube, took at 3000x5msec cube at max speed -> cube 0032

09:36 pauseGuiding

09:37 +0.8 mm 300x80ms@10Hz cube 0033

09:39 -0.8 mm 300x80ms@10Hz cube 0034

Open the side vent doors. Rear were already open.

09:42 still at -0.8 mm 1000x80msec at max speed just watching the seeing go by cube 0035

09:46 background without star 300x80ms@10Hz cube 0036

Data to measure the wavefront drift in open loop collimation

mkdir /RepositoryAGW_Data/20081019w/

09:50 pointto M5_0269 gs=7 Tel="ACTIVE"

10:00 stopWFS to end GCS active optics

10:05 agw_collimate in IDL (with funny interlude teach Andrew to spell IDL in Linux - we did have a case mistake in the instructions!)

10:09 agw_wfsc_multiple, 40, EXP=30 starting with wfsc 0006

10:15 Andrew starts adjusting focus (pressing Collimate after each exposure starting now)

10:21 Star starts bouncing left-right on IRTC

10:26 Experiment ended by cloud. Guider drives star off the field because of the 100-count gradient across the image.

Half hour pause for clouds (tea break)

11:06 Back at work since clouds are gone.

Johannes says seeing is 0.9 arcsec.

Looking for focus hysteresis

11:08 pointo BS9122 gs=0 Tel="GUIDE" at EL=68

11:11 Clear C00

11:12 agw_collimate starting with wfsc 0027-32

11:18 driving in +2000 Z4 on Left M1 wfsc 33-35

11:21 driving z4 to +1000 from initial position (reducing by 1000 from previous position). wfsc 36-38

11:25 driving Z4 back to nominal 0 (reducing by 1000 from previous measurement). wfsc 39-41

11:27 driving focus to -1000 nmn wfsc 42-44

11:31 driving focus to -2000 wfsc 45-47

No focus hysteresis but SCALE FACTOR is wrong by a factor of 2!

Checking Scale Factors on the WFS ZErnike Terms

11:35 now check other coefficients, first collimate... wfsc 48-49

11:37 driving in 1000 nm Z5 wfsc 50-52

11:40 now going to -1000 Z5 wfsc 53-55

11:43 now collimating wfsc 56

11:44 now +1000 Z6 wfsc 57-59

11:47 now -1000 Z6 (-1000 from initial position) wfsc 60-62

11:49 now collimating wfsc 63

11:52 now sending +1000 nm Z7 wfsc 64-66

11:54 now Z7 to -1000 nm wfsc 67-69

11:56 now collimate wfsc 70

11:57 now Z8 to 1000 nm wfsc 71-73

12:00 now Z8 to -1000 wfsc 74-76

12:03 now collimate wfsc 77-79

12:06 now + 1000 nm Z9 wfsc 80-82

12:09 now -1000 nm Z9 wfsc 83-85

12:11 now collimate wfsc 86

12:13 now +1000 Z10 wfsc 87-89

12:15 now -1000 Z10 wfsc 90-92

12:18 now collimate wfsc 93

12:20 now +1000 z11 wfsc 94-96

12:22 now - 1000 z11 wfsc 97-99

NOTE THE BIG FOCUS GOING WITH Z11

12:25 collimate on axis wfsc 100

12:26 now send +500 Z22 but get "bad forces" error, C00 and try again

12:27 collimate wfsc 101-104

12:30 now try again +500 nm Z22 still bad forces, have ascertained that I can go -500nm so will C00 collimate and try again... (Maybe PSF isn't forgetting the previous Z22 request when the bad forces error happens?)

12:35 collimate wfsc 105-108

12:37 -500 nm Z22 wfsc 109-111

12:38 collimate wfsc 112

12:39 send -500 nm Z22 again but this time with the "FCin" flag set to 0 so that focus moves with M1 Z motion. wfsc 113-115 It looks the the bending was applied without the focus term, but the Z-active optics position wasn't adjusted.

12:41 collimate wfsc 116

12:43 quadrafoil (Z14) +500 nm wfsc 117-119

12:44 quadrafoil (Z14) -500 nm wfsc 120-122

12:45 collimate wfsc 123

12:46 quadrafoil (Z15) +500 nm wfsc 124-126 noticing coupling between Z14-15 so I will now back out the 500 nm Z15 and try another set of measurements to check...

sun is up and tadpoles come out to play...

13:08 Close Dome

-- JohnHill - 19 Oct 2008
Topic revision: r8 - 28 May 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback